OpenStudy (johnnydicamillo):

Find the solution of the logarithmic equation logbase(2)(x^2-4x-28)=2

OpenStudy (johnnydicamillo):

\[\log _{2}(x^2-4x-28)=2\]

OpenStudy (paxpolaris):

take 2^ both sides

OpenStudy (johnnydicamillo):

what do you mean?

OpenStudy (paxpolaris):

\[\Large 2^{\log _{2}(x^2-4x-28)}=2^2\]

OpenStudy (paxpolaris):

on the left "2 to the power of" cancels logbase2... \[x^2-4x-28=4\]

OpenStudy (johnnydicamillo):

okay, so I need to factor correct?

OpenStudy (paxpolaris):

yes

OpenStudy (johnnydicamillo):

okay so (x+4) (x-8) = 0?

OpenStudy (paxpolaris):

right. x=-4 or x=8 both answers work if you plug in the original equation.